Give the empirical formula of each of the following compounds if a sample contains the following quantities of each element.


(A) 0.0334 mol C, 0.1000 mol H, and 0.0167 mol O


(I have tried C2H3O and it was incorrect)


1
Expert's answer
2021-02-25T03:12:45-0500

(a) 0.0334 mol C, 0.1000 mol H and 0.0167 mol O

C = "\dfrac{0.0334}{0.1000}=0.334"

H = "\dfrac{0.1000}{0.1000}=1"

O = "\dfrac{0.0167}{0.1000}=0.167"

Multiplying by suitable integer 6

"C=0.334\times6=2.004\approx2"

"H=1\times6=6"

"O=0.167\times6=1.002\approx1"

Empirical Formula "=C_2H_6O"